OpenStudy (anonymous):

Rhett decides to build a square room for his movie and music collection. If the area of the room is 4x^2 + 28x + 49 square feet, what is the length of one side of the room?

Directrix (directrix):

Area of Square = side * side 4x^2 + 28x + 49 = (side)^2 (2x + 7) ^2 = (side) ^2 2x + 7 = side of room